Pupil Check after OCAm fix during the SSD.

The OCAM was fix to the bench, this activity was done in the lab during the summer shutdown.

Alessandro and Juan Carlos worked after the OCAM fix to measure alignments parameters such as the interdistance vs modulation gives a relationship to the focus.

The AGW is mounted at the telescope and initail check of the pupil interdistance shows the same values than in the lab.

Waiting to have the Adsec installed and the option to close the loop with a few modes and low gain to allow a measurements of the pupil wobble.

The measurements were span in three days before the flatteinign task.

The measurements are taken with the RR configuration.

The tasks were diveded into:

1.- identify and center the return beam.

2.- initial estimation of the pupil shift and execute a full rotation at Bin1.

3.- Check the Bin4 pupil position. First estimation of the pupils. Rotations test.

4.- interdistance with the RR.

A log of the activies is found in:

GOOGLE DRIVE LOG.

1.- Return spot found. Tbis process took about 3 hours, this is a long process, it required the spiral search.

The spot moves from one day to another, so it requires a small scann in the Rx and Ry to get the return back.

2.- The ocam looks like well centered anf using the origila pupil for Bin1, the wobble is inside of the pupil wobble. Not shift required.

The puils has been tested in close loop and it performed well.

image of pupil wobble.

bin1_20220830.png

3.- Bin4 using the original pupils shows to be just at the limit with a saturation of the camera lens in the Y negative, with a pupil error <=0.3.. The position will allow a movement in negitve direction ofr Y. The pupil shifted shows to be in range for the Y axis.

Image before and after the shift.

bin4_before-aftershift_20220830.png

4.- Interdistance measurement.

This is a very tricky procedure that it is affected of how weel the system is correcting, vibrations etc.

Discussed with Enrico, Simone and Doug (Night 30th), collected data for posprocessing. The values useful are from modulation 20 to 40l/D. Enrico will take the TN and posprocessing the data to get an estimation.

A part of the pupil check and interdistance, it was performed a close loop with Bin1. The night 30th doug, reported serius instabilitues of the close loop.

In a series of close loop using the 10 new acurators and added and removed from the actuator list showed a good performences and a slightly improvement with the 10 actuator back in operations. Data were taken for postprocessing.

Conclusion.

From Simone and Enrico comments, the OCAM should be leave as it is. to try to move a small amout will bring another problem. The pupil wobble is in rage of the CL.

About the interdistance , data were taken for Enrico post processing, but in a first approach of the data, the system is not with one pixel out of focus.

The close loop instability not shown, and it conclude that the use of 10 in/out for the close loop give a small gain, but the loop is stable with both configurations.

Coming tasks.

  • Check pupil stabilities.
  • Verify the pupil for Bin4.
  • Reconstruct Bin2 and Bin3 pupils and check.

-- JuanCarlosGuerra - 31 Aug 2022
Topic revision: r2 - 30 Sep 2022, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback